A heat engine operates in a Carnot cycle between 86.0°C and 345°C. It absorbs 20,000 J of energy per cycle from the hot reservoir. The duration of each cycle is 5.00 s.

(a) What is the mechanical power output of this engine?

(b) How much energy does it expel in each cycle by heat?
Answer #1

Ane - Cliven Tez 86°C = ( 273 +86) k 2 359k TH=3450 2 (273 +345) K = 618k Bu dabi Qab = 200007 ta 5 see. Efficiency of heat e